OpenStudy (anonymous):

33% is to 66 as (x) is to 15 what is x

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

x is 7.5 % See 33% = 33/100 and 33/100 x 200 = 66 so for 15 7.5/100 x 200 = 15

OpenStudy (anonymous):

if u r satisfied, pls click the GOOD ANSWER button to mark the question as solved and closed